OpenStudy (anonymous):

Integral of (x-2)^(-1/3) from 1 to 3. I get 0. But Wolfram gets imaginary numbers!!!!! Why is that? Who is right?

OpenStudy (ipwnbunnies):

Nah, you cray. It's a simple power rule integral.

OpenStudy (ipwnbunnies):

\[\int\limits_{1}^{3} (x-2)^{-1/3} dx\]

OpenStudy (ipwnbunnies):

u = (x-2); du = 1 dx \[\int\limits_{1}^{3} u^{-1/3} du\]

OpenStudy (ipwnbunnies):

Can you take it from here?

OpenStudy (anonymous):

No. This is improper integral.

OpenStudy (anonymous):

FTC does not apply.

OpenStudy (whpalmer4):

Need to change the limits of integration when you change variables but the indefinite integral will be \[\frac{3}{2}(x-2)^{2/3} + C\] The problem is evaluating \(-1^{2/3}\)

OpenStudy (anonymous):

\[\int\limits_{1}^{3}\left( x-2 \right)^{\frac{ -1 }{ 3 }}dx=\frac{ \left( x-2 \right)^{\frac{ -1 }{ 3 }+1 } }{ \frac{ -1 }{ 3 }+1 }~x~ from~ 1 \to 3\] \[=\frac{ 3 }{ 2 }\left[ \left( 3-2 \right)^{\frac{ 2 }{ 3 }} -\left( 1-2 \right)^{\frac{ 2 }{ 3 }}\right]\] \[=\frac{ 3 }{ 2 }\left[ 1^{\frac{ 2 }{ 3 }} -\left( \iota^2 \right)^{\frac{ 2 }{ 3}}\right]\] \[=\frac{ 3 }{ 2 }-\frac{ 3 }{ 2 }\iota ^{\frac{ 4 }{ 3 }},\] which is an imaginary number.

OpenStudy (anonymous):

I think I should consider i^(4/3) to be 1 simply.

OpenStudy (whpalmer4):

\[\int_1^3 \frac{1}{\sqrt[3]{x-2}} \, dx = -\frac{3}{2}(1+(-1)^{2/3})\] Since when does \(i^{4/3} =1\)?

OpenStudy (anonymous):

\[cube~roots~of~1~are 1,\omega,\omega^2 \] \[1^4=1,\omega^4=\omega^3\omega=1\omega=\omega,\left( \omega^2 \right)^4=\omega^8=\left( \omega^3 \right)^2\omega^2=1^2\omega^2=\omega^2\] so you can not consider it as 1.

OpenStudy (anonymous):

When using WolframAlpha you have to specify that you're taking the real root. It's designed to find the principal root by default.